# Alignment-Aware Model Extraction Attacks on Large Language Models

### v1

@@ C.2 Proofs of Theorem 1
Guarantee of KD. As we know, $\mathbb{D}_{KL}(p,q)\geq 0~{}\forall~{}p~{}\text{and}~{}q$. Therefore, if $\mathcal{L}_{kd}$ shown in Equation 14 equals to 0, then both $\mathbb{D}_{KL}[P_{\theta}(\cdot|\mathbf{x})||P_{\theta_{vic}}(\cdot|\mathbf{x})]$ and $\mathbb{D}_{KL}[\text{SM}(P_{\theta}(\cdot|\mathbf{x})/T)||\text{SM}(P_{\theta_{vic}}(\cdot|\mathbf{x})/T)]$ equal to 0. For the latter one, we have
$\mathbb{D}_{KL}[\text{SM}(P_{\theta_{vic}}(\cdot|\mathbf{x})/T)||\text{SM}(P_{\theta}(\cdot|\mathbf{x})/T)]$
$=\mathbb{E}_{\mathbf{y}\sim P_{\theta_{vic}}(\cdot|\mathbf{x})}\mathbb{E}_{y\in\mathbf{y}}$
${\text{log}\frac{\text{exp}({P}_{\theta}(y|\mathbf{x},\mathbf{y}_{p})/T)/\sum_{y^{\prime}\in\mathbf{y}}{\text{exp}({P}_{\theta_{vic}}(y|\mathbf{x},\mathbf{y}_{p})/T)}}{\text{exp}({P}_{\theta_{vic}}(y|\mathbf{x},\mathbf{y}_{p})/T)/\sum_{y^{\prime}\in\mathbf{y}}{\text{exp}({P}_{\theta}(y|\mathbf{x},\mathbf{y}_{p})/T)}}}$
$=\mathbb{E}_{\mathbf{y}\sim P_{\theta_{vic}}(\cdot|\mathbf{x})}\mathbb{E}_{y\in\mathbf{y}}$
${\text{log}\frac{\text{exp}(({P}_{\theta}(y|\mathbf{x},\mathbf{y}_{p})-{P}_{\theta_{vic}}(y|\mathbf{x},\mathbf{y}_{p}))/T)}{\sum_{y^{\prime}\in\mathbf{y}}{\text{exp}({P}_{\theta_{vic}}(y|\mathbf{x},\mathbf{y}_{p})/T)}/\sum_{y^{\prime}\in\mathbf{y}}{\text{exp}({P}_{\theta}(y|\mathbf{x},\mathbf{y}_{p})/T)}}},$
where we can observe that only when $P_{\theta}(\cdot|\mathbf{x})$ equals to $P_{\theta_{vic}}(\cdot|\mathbf{x})$ can this term reduce to 0. Integrating the analysis of these two terms, we can obtain that $\mathcal{L}_{kd}=0$ represents the local model’s distribution converge to that of the victim model.
Guarantee of LoRD. When $\mathcal{L}$ shown in Equation 12 equals to 0, the proportion of $P_{\theta_{t}}(\mathbf{y}_{{vic}}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{-}|\mathbf{x})$ and $P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{+}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{-}|\mathbf{x})$ should limit to $\infty$- ∞. As we know that i) in a distribution $\sum{P_{\theta_{t}}(\cdot|\mathbf{x})}=1$ and ii) $\mathbf{y}^{+}_{t-1}$ is a dynamic positive response generated at each period, we can deduct that when $\mathcal{L}=0$ there must be $\mathbf{y}_{vic}=\mathbf{y}_{t-1}^{+}$, i.e., $P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})=P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})=1$ and $P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})=0$. Note that this is merely a theoretical limit that cannot be reached, because $\mathbf{y}_{t-1}^{-}$ will not be sampled if its probability is 0, and $\mathbf{y}_{t-1}^{+}$ usually doesn’t exhibit a significant distinction to $\mathbf{y}_{t-1}^{-}$ when sampling.

### v2

@@ B.3 The Proof of Proposition 2
Guarantee of KD. As we know, $\mathbb{D}_{KL}(p,q)\geq 0~{}\forall~{}p~{}\text{and}~{}q$. Therefore, if $\mathcal{L}_{kd}$ shown in Equation 14 equals to 0, then both $\mathbb{D}_{KL}[P_{\theta}(\cdot|\mathbf{x})||P_{\theta_{vic}}(\cdot|\mathbf{x})]$ and $\mathbb{D}_{KL}[\text{SM}(P_{\theta}(\cdot|\mathbf{x})/T)||\text{SM}(P_{\theta_{vic}}(\cdot|\mathbf{x})/T)]$ equal to 0. For the latter one, we can derive that only when $P_{\theta}(\cdot|\mathbf{x})$ equals to $P_{\theta_{vic}}(\cdot|\mathbf{x})$ can this term reduce to 0 based on the property of KL divergence. Integrating the analysis of these two terms, we can obtain that $\mathcal{L}_{kd}=0$ represents the local model’s distribution converge to that of the victim model.
@@
Guarantee of LoRD. When $\mathcal{L}$ shown in Equation 11 equals to 0, the proportion of $P_{\theta_{t}}(\mathbf{y}_{{vic}}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{-}|\mathbf{x})$ and $P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{+}|\mathbf{x})/P_{\theta_{t}}(\mathbf{y}_{{t-1}}^{-}|\mathbf{x})$ should limit to $\infty$- ∞. As we know that i) in a distribution $\sum{P_{\theta_{t}}(\cdot|\mathbf{x})}=1$ and ii) $\mathbf{y}^{+}_{t-1}$ is a dynamic positive response generated at each period, we can deduct that when $\mathcal{L}=0$ there must be $\mathbf{y}_{vic}=\mathbf{y}_{t-1}^{+}$, i.e., $P_{\theta_{t}}(\mathbf{y}_{vic}|\mathbf{x})=P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})=1$ and $P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})=0$.
Note that this is merely a theoretical limit for LoRD that cannot be reached, because $\mathbf{y}_{t-1}^{-}$ will not be sampled if its probability is 0, and $\mathbf{y}_{t-1}^{+}$ usually doesn’t exhibit a significant distinction to $\mathbf{y}_{t-1}^{-}$ when sampling.
Appendix C Supplemental Related Works
C.1 Human-Feedback-Free Alignments
There are several alternatives to the standard RLHF approach. Lee et al. (2023) propose reinforcement learning with AI feedback (RLAIF) as a means to diminish the annotation burden associated with the preference assessments. Besides, there are some approaches, such as direct preference optimization (DPO) (Rafailov et al., 2023), that conceptualize the language model itself as the reward model and thus consolidate Equation 5 and Equation 6 into a unified supervised and preference-based training task. Since they do not change the primary targets (i.e., maximizing rewards) and optimization strategies of LLM’s alignments, we only consider the standard formation of alignments for simplicity in our theoretical analysis.
C.2 Language Models Extraction
Table 9: Datasets and pre-trained model checkpoints used in the paper. Specifically, we select twelve representative datasets: SafeRLHF Ji et al. (2024), DiaSafety Sun et al. (2022), WMT16 Bojar et al. (2016), TLDR Kirk et al. (2023), CNN Daily Mail Hermann et al. (2015), Samsum Gliwa et al. (2019), WikiSQL Zhong et al. (2017), Spider Yu et al. (2018), E2E-NLG Dušek et al. (2020), CommonGen Lin et al. (2020), PIQA Bisk et al. (2020), and TruthfulQA Lin et al. (2021) as benchmarks for our evaluation. These datasets cover most of the downstream tasks in natural language generation.
Studies to steal language models originated from the natural language understanding (NLU) models, such as BERT(Devlin et al., 2019), and then evolved to generative language models, especially large language models recently.
Krishna et al. (2020) highlights early recognition of model extraction threats in language models. By constructing text inputs with randomly vocabulary sampling, they successfully extract the weights from BERT-based APIs. Besides, Rafi et al. (2022) investigate the feasibility of side-channel model extraction attacks, revealing that by analyzing extra signals from GPU kernels, one could accurately steal the model architecture and its parameters. Subsequent research (Xu et al., 2022) has thoroughly investigated the strategy of ensembling victim models to train a competitor model that surpasses its teachers.
The exploration of generative language model extraction is still in its infant stage, with only a handful of studies thus far. Wallace et al. (2020) investigate imitation attacks on natural language models. By designing monolingual query texts and collecting responses, they successfully extract the knowledge from a simulated machine translation model under the black-box settings. This research exhibits that slight architectural differences will not influence the extraction between language models. Li et al. (2023b) also explores the potential risks of stealing the code-generation abilities of LLMs into smaller downstream models. Unlike previous research (Wallace et al., 2020), this is the first study that selects LLMs as targets. By collecting large-scale domain-specific samples, they fine-tune a 7-billion local pre-trained model with them and show the similarity between the victim and local models in both performances and adversarial samples. However, these two studies employ the MLE loss (Equation 3) as the MEA method, neither considering whether MLE is compatible with LLMs’s training, especially the alignment procedure shown in Section 2.2, nor addressing optimizations related to query efficiency and the watermark resistance. Besides, the scope of these studies is limited to stealing specific knowledge in a few downstream domains. At the same time, most of the critical aspects of LLMs and the required extraction capabilities, such as query numbers and local model scales, remain unresolved.
C.3 Text Watermarks
In contrast to stealing LLMs, IP protection methods have received considerable attention recently. By sampling a stealthy but representative “greed word set” on the vocabulary distribution, these methods (Cong et al., 2022; He et al., 2022, 2021; Kirchenbauer et al., 2023) can remap the generated words into their synonyms or add the “watermarked” token automatically, and thus effectively certify the output. Besides, strategies such as integrating embeddings into the representation as the backdoor (Peng et al., 2023) or manipulating the probabilities with crafted sinusoidal noises (Zhao et al., 2022, 2023) are also proposed. However, these approaches often presume more stringent conditions regarding the victim and the suspected models. This paper will further assess the effectiveness of LoRD and current MEAs in evading these black-box watermarking strategies.
Appendix D A Detailed Threat Model
Adversary’s Objective. The adversary’s objective is to steal the targeted knowledge from LLMs. Specifically, we select machine translation, reasoning, data-to-text, structured text generation, and summarization as the downstream domain-specific tasks. The adversary aims to develop a query-efficient MEA algorithm, since the amount of input and generated tokens will be counted as the costs. Additionally, the MEA methods are expected to be watermark-resistant, i.e., they are highly desired to reduce the risks of exposure to unauthorized stealing.
Algorithm 1 LoRD Algorithm
1:  Input:Query dataset $\mathcal{D}_{q}$, local language model $\theta_{init}$, interface of the victim model $P_{\theta_{vic}}(\cdot|\cdot)$, train period number $N_{t}$, threshold values $\tau_{1}$ and $\tau_{2}$.
2:  // Initialization.
3:  $\theta_{0}\leftarrow\theta_{init},\mathcal{D}_{tr}\leftarrow\emptyset,\mathcal{D}_{0}^{+}\leftarrow\emptyset,\mathcal{D}_{0}^{-}\leftarrow\emptyset,t\leftarrow 0$;
4:  // Query the victim model.
5:  for $\mathbf{x}\sim\mathcal{D}_{q}$ do
6:     $\mathbf{y}_{vic}\leftarrow P_{\theta_{vic}}(\cdot|\mathbf{x})$;
7:     $\mathcal{D}_{tr}\leftarrow\mathcal{D}_{tr}\cup\{(\mathbf{x},~{}\mathbf{y}_{vic},~{}P_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x}))\}$;
8:  end for
9:  // Train local model.
10:  // Initialize the positive and negative datasets.
11:  $\mathcal{D}_{0}^{+}\leftarrow\mathcal{D}_{q}$;
12:  for ($\mathbf{x},\mathbf{y}_{vic},{P}_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x}))\sim\mathcal{D}_{tr}$ do
13:     $\mathbf{y}_{0}^{-}\sim P_{\theta_{t}}(\cdot|\mathbf{x})$;
14:     $\mathcal{D}_{0}^{-}\leftarrow\mathcal{D}_{0}^{-}\cup\{(\mathbf{x},\mathbf{y}_{0}^{-},P_{\theta_{0}}(\mathbf{y}_{0}^{-}|\mathbf{x}))\}$;
15:  end for
16:  while $t<N_{t}$ do
17:     $t\leftarrow t+1$;
18:     $\theta_{t}\leftarrow\theta_{t-1}$;
19:     for ($\mathbf{x},\mathbf{y}_{vic},{P}_{\theta_{vic}}(\mathbf{y}_{vic}|\mathbf{x}))\sim\mathcal{D}_{tr}$ do
20:        $\mathbf{y}_{t}^{+},\mathbf{y}_{t}^{-}\sim P_{\theta_{t}}(\cdot|\mathbf{x})$;
21:        $\mathcal{D}_{t}^{+}\leftarrow\mathcal{D}_{t}^{+}\cup\{(\mathbf{x},\mathbf{y}_{t}^{+})\}$;
22:        $\mathcal{D}_{t}^{-}\leftarrow\mathcal{D}_{t}^{-}\cup\{(\mathbf{x},\mathbf{y}_{t}^{-})\}$;
23:     end for
24:     // Forward.
25:     for $\mathbf{x},\mathbf{y}_{vic},\mathbf{y}_{t-1}^{+},\mathbf{y}_{t-1}^{-}\sim(\mathcal{D}_{tr},\mathcal{D}_{t-1}^{+},\mathcal{D}_{t-1}^{-})$ do
26:        $\Delta^{+}\leftarrow\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|x)-\text{log}P_{\theta_{t-1}}(\mathbf{y}_{t-1}^{+}|x)$;
27:        $\Delta^{-}\leftarrow\text{log}P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|x)-\text{log}P_{\theta_{t-1}}(\mathbf{y}_{t-1}^{-}|x)$;
28:        if  $\Delta^{+}<\Delta^{-}$ then
29:           swap$(\mathbf{y}_{t-1}^{+},\mathbf{y}_{t-1}^{-})$;
30:           swap$(\Delta^{+},\Delta^{-})$;
31:        end if
32:        if  $P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|x)<\tau_{1}$& $\Delta^{+}<\tau_{2}$ then
33:           $\mathbf{y}_{t-1}^{+}\leftarrow\mathbf{y}_{vic}$;
34:        end if
35:        // Compute loss with Equation 10 or 11.
36:        $\mathcal{L}\leftarrow\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{t-1}^{+}|\mathbf{x})}]+clip(\text{log}[\frac{P_{\theta_{t}}(\mathbf{y}_{t-1}^{-}|\mathbf{x})}{P_{\theta_{t}}(\mathbf{y}_{\text{vic}}|\mathbf{x})})]$;
37:        // Backward.
38:        $\theta_{t}\leftarrow\text{stepUpdate}(\theta_{t},\mathcal{L})$;
39:     end for
40:  end while
41:  return  $\theta_{t}$
Targeted Models. We select Llama3-70B, GPT-3.5-turbo, and GPT-4o as the victim models in this paper. Unlike previous works that only deployed simulated local victim models (e.g., OPT (Zhang et al., 2022)), our selections aim to expose the stealing threat on realistic AI services. Besides, our target models are specifically constrained to LLMs fine-tuned with alignment methods (e.g., RLHF) since they are not only state-of-the-art solutions now but also more valuable due to their human-based alignments.
Adversary’s Capabilities. In accordance with the LLM-based AI service APIs, we identify two attack scenarios: black-box and grey-box attacks. In the black-box scenario, only textual responses the adversary is allowed to obtain. At the same time, all other information, such as the temperature, sampling strategies, and the hidden states of LLMs, are unseen and inaccessible. On the contrary, a grey-box attack allows the adversary to access the generation probabilities distribution of tokens. Notice that both MLE and our LoRD method are under black-box settings, and we only adopt grey-box settings on some particular stealing methods, such as knowledge distillation.
Besides, this paper posits that the adversary usually has worse training conditions than the victims. Specifically, query times and the scale of the local model available to the adversary are much smaller than the victims’ training datasets and model parameters. This setting has been adopted in previous LLMs’ extraction (Li et al., 2023b). We call it a LaViSH (Large-Victim-Small-Heist) framework, which allows us to estimate the upper bound of MEA risks empirically. For adversaries with more substantial resources, they can train more powerful MEA-based LLMs by leveraging MEA algorithms under our LaViSH settings.